### v1

Abstract
Vertical Federated Learning (VFL) is an emergent distributed machine learning paradigm wherein owners of disjoint features of a common set of entities collaborate to learn a global model without sharing data. In VFL, a host client owns data labels for each entity and learns a final representation based on intermediate local representations from all guest clients. Therefore, the host is a single point of failure and label feedback can be used by malicious guest clients to infer private features. Requiring all participants to remain active and trustworthy throughout the entire training process is generally impractical and altogether infeasible outside of controlled environments. We propose Decoupled VFL (DVFL), a blockwise learning approach to VFL. By training each model on its own objective, DVFL allows for decentralized aggregation and isolation between feature learning and label supervision. With these properties, DVFL is fault tolerant and secure. We implement DVFL to train split neural networks and show that model performance is comparable to VFL on a variety of classification datasets.
1 Introduction
Federated Learning McMahan et al. (2017), or FL, was introduced by Google researchers as a strategy for distributed learning, addressing communication efficiency and data privacy. Distributed participants in FL training do not expose their data to any other party. Instead, they train local models on their data, guided by a coordinator that has no knowledge of the agents’ data. Thus, FL extends collaborative machine learning to contexts where data communication is undesirable or forbidden. Recent advances in communication systems and cloud computing infrastructure combined with an increased public awareness and legislation regarding data privacy (such as the EU’s GDPR) have contributed to FL’s advancement as an emergent machine learning paradigm.
Among FL’s developments is Vertical Federated Learning (VFL). Standard FL can be said to have a “horizontal” or “sample-parallel” division of data, i.e. each participant holds unique samples within a shared feature space. On the other hand, VFL participants hold unique features of a common sample space, which is a “vertical” division, and is “feature-parallel”. This allows guest agents who individually have incomplete information about their target to learn meaningful joint representations without sharing data. A typical example: an imaging center, pathology lab and general hospital may come together to predict rare types of cancer, even though health regulations such as HIPAA may disallow them from collecting each others’ data.
VFL guest agents undertake an “entity alignment” phase before training, where they ensure their datasets’ horizontal indexing is identical, i.e., all participants process features belonging to the same entity in any round of training. At each round, each guest passes its features to its local “guest model.” Guests share their local representations with a host. For example, a guest training a vertically federated ANN may send its activations. A boosted decision tree guest may send a weak learner. The host owns data labels and is aware of the identity of the sample or batch that the guests are processing at any time. The host also possesses a “host model” capable of aggregating these representations and making predictions against the labels. The host calculates a loss based on its aggregation and provides feedback (e.g. gradients) to the guests, which then begin the next round of training. We refer to the combined system of guest and host models as VFL models.
VFL has found increasing success in recent years, with various real-world implementations and proposals Liu et al. (2023). However, VFL suffers from some flaws which make its implementation impractical where participants are few in number and each computationally powerful (as in cross-silo scenarios), and entirely infeasible otherwise. One such flaw is that the link between the guests and the host is a single point of failure. Training a batch of data is contingent on the exchange of intermediate results between the guests and the host. If the host waits indefinitely for all guests to communicate, a guest crashing or losing connection would completely halt training. If the host sets a deadline, some inputs to its model will remain unfulfilled. (In the latter case, a late message will also cause a catastrophic fault). If gradients from the host to the guests are lost or arrive in the incorrect order, guest model convergence is hampered. One might dismiss such events as infrequent or at least controllable in cross-silo settings. However, that is naive. When some of the guests (and even the host) could be geographically sparse and/or low-power devices, it is catastrophic.
A fundamental limitation of VFL is imposed by the size of the joint sample space. For instance, databases in groups of large nationwide financial institutions may share many but not all customers. Guests must discard data outside the intersection, which could have been used to learn better guest models. In systems where the intersection is small, there may not be enough samples to train a VFL model altogether. In general, this also hinders VFL’s applicability to systems with large-scale participation. Consider again the hospital example. Suppose the task is now to predict bone marrow cancer. Along with the pathology lab, the imaging center and the general hospital, we also include data collected by an orthopaedic hospital. Even if the overlap of patients at the other facilities is significant, their overlap with patients of the orthopaedic hospital may be too small.
Another drawback of VFL (and FL in general) is the risk of inference attacks from curious guests. Intuitively, a host makes predictions based on information it receives from all guests. The updates or feedback it returns is derived from this prediction and the ground truth. Therefore, feedback sent to one guest contains some mutual information with respect to the others and to the label. Luo et al. (2021); Fu et al. (2022) show that a malicious agent can engineer its messages to the host such that the label and the other guests’ features can be conveniently extracted from the gradients it receives in return. Attempts to safeguard VFL against inference attacks are generally based on differential privacy, homomorphic encryption, and multi-party techniques such as audits and shared secrets Yang et al. (2023); Khan et al. (2023). These operations incur a high computational cost, which might not be feasible outside some cross-silo scenarios. Further, these methods hamper model convergence.
Building intricate mechanisms to mitigate these issues inflates the complexity of the problem. In this work, we attack the source directly. Backpropagating gradients to guests allows inference attacks. The wait for a response or feedback causes synchronization issues. What if we eschewed end-to-end backpropagation (BP) altogether? We present Decoupled Vertical Federated Learning (DVFL), a novel strategy for ANNs to address these shortcomings. The underlying principle is blockwise learning, where gradient-isolated groups of layers of an ANN are trained on local objectives. Such greedily trained ANNs have been shown to perform comparably to those trained by BP Belilovsky et al. (2019); Journé et al. (2023). We decouple guest models from host models by training them on their own local objectives. Similarly, we decouple supervision from aggregation by training host models on local objectives. We use a final transfer learning model to make predictions against labels. This yields two useful system properties. One, this allows redundant hosts. Standard VFL would require hosts to share labels with each other, which is forbidden. Two, we may use separate datasets for feature extraction and label inference.
With redundancy, even if a guest’s connection to one host fails, training may continue as normal on others. If hosts may fail, the redundancy also affords some degree of tolerance. By using separate datasets for inference and representation learning, we may train guest and host models on data which is not necessarily labeled and not necessarily a part of a joint sample space. This makes DVFL applicable to systems with small intersections of sample spaces. Further, the label owner’s supervised model can be trained completely offline, without the risk of faults.
In Section 2, we provide a formal description of VFL and some algorithms which implement it for various machine learning models. We also elaborate on the shortcomings we hope to address with DVFL. In Section 3, we review some closely-related lines of research and use them to motivate DVFL. In Section 4, we provide a formal description of DVFL with an implementation for ANNs. We show some baseline results on benchmark machine learning tasks in Section 5. In the Appendix, we present more results and discuss our design choices. The main contributions of this work are:
  1. 1.
    We present DVFL, a strategy which eschews host-to-guest feedback as well as label-to-host feedback and therefore has inbuilt redundancy, security and fault handling (including faults during the backward pass) while performing comparably to VFL.
  2. 2.
    We show host redundancy improves model performance, especially when there are faults.
  3. 3.
    We demonstrate how DVFL guests might exploit (possibly unlabeled) data outside a limited sample intersection to learn strong representations.
2 Background: Standard VFL
Entities $\mathbf{x}$ are realizations of a fixed distribution $X$ (say). The rows of a dataset are features of a unique entity $\mathbf{x}$. We may serialize these entities using a sample index. Let us say the set of sample indices corresponding to entities the host has labels for is $\mathcal{S}_{\mathrm{labels}}$. Let us say the entity with index $j$ is $\mathbf{x}_{j}$.
2.1 The VFL Hierarchy
Guests: $\mathcal{G}$ is the set of feature-owning collaborators: the guests. ||$\cdot$|| ⋅ | denotes cardinality. Each $g_{i}\in\mathcal{G}$ has its own
  1. 1.
    Dataset $\mathcal{D}_{i}=\left\{\mathbf{x}_{j,\mathcal{F}_{i}}\in\mathbb{R}^{|\mathcal{F}_{i}|}:j\in\mathcal{S}_{i}\right\}$, where $\mathcal{F}_{i}\subseteq\{1,2,\cdots,D\}$ is the subset of $D$ features that the guest owns, and $\mathcal{S}_{i}$ is the set of entities $g_{i}$’s dataset has records for. $\mathbf{x}_{j,\mathcal{F}_{i}}$ are the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$.
  2. 2.
    Guest model $m_{i}(\text{ }\cdot\text{ };\theta_{i}):\mathbb{R}^{|\mathcal{F}_{i}|}\to\mathbb{R}^{k_{i}}$ with parameters $\theta_{i}$.
The guest model $m_{i}$’s task is to concisely represent the features $\mathcal{F}_{i}$ of $\mathbf{x}_{j}$, for all sample indices $j$ that are in the intersection of all guests, i.e. $\forall j\in\mathcal{S}_{\mathrm{guests}}$, where $\mathcal{S}_{\mathrm{guests}}:=\bigcap_{i=1}^{|\mathcal{G}|}\mathcal{S}_{i}$ and pass them to the host $h$. Guests do not have access to other guests’ model parameters or data features. In general, they do not have access to labels, either.
Host: The host agent $h$ executes every round of training. The host has access to target labels. The host, given representations for all available features of $\mathbf{x}_{j}$, may use its own model to make useful predictions for $\mathbf{x}_{j}$. The host has
  1. 1.
    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
  2. 2.
    Host model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{h}$.
It is possible that $h\in\mathcal{G}$ i.e. the host may also be one of the guests. In such situations, the host is often known as the “active party” and guests as “passive parties.”
If the host has access to the full computational graph, it may adjust all parameters to better predict label $\mathbf{y}$, usually by minimizing the expected value of some loss function. Otherwise, it may update only its parameters $\theta_{h}$ and send intermediate results (e.g. gradients) via which the guests may update their own models. Together, the guests and host solve the following joint optimization problem:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}_{j}\sim X}\left[\ell(m_{h}\left(\hat{\mathbf{x}}_{j,1},\dots,\hat{\mathbf{x}}_{j,{|\mathcal{G}|}});\theta_{h}\right),\mathbf{y}_{j})\right]\,,$
where $\hat{\mathbf{x}}_{j,i}\equiv m_{i}(\mathbf{x}_{j,\mathcal{F}_{i}};\theta_{i})\in\mathbb{R}^{k_{i}}$.
2.2 Private Set Intersection & Entity Alignment
All guests process entities in the same order, each passing the next entity to its local model. For this to happen, all guests must know the set of common entities $\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$. This is usually achieved using entity resolution (if required) and some form of Private Set Intersection (PSI).
PSI is a multiparty computation of the intersection of sets belonging to each party without revealing members outside the intersection, typically by exchanging encrypted versions of the set Freedman et al. (2004). Entity resolution is a multiparty computation to generate a set of unique identifiers from features which are not strictly unique or precise. Entity resolution may be performed by comparing cryptographic long keys Hardy et al. (2017); Nock et al. (2018). After the host obtains the results of PSI, there is a phase of entity alignment. In this phase, the host decides the order of entities which all guests must follow and informs them.
2.3 Split Training for Neural Networks
A faithful implementation of VFL for ANNs is SplitNN Gupta & Raskar (2018). SplitNN was originally proposed as an algorithm to distribute the training of a global neural network across two agents by splitting the network depth-wise into two. At the “split layer”, forward and backward signals are communicated between the two devices. [^1] One might say that the standard SplitNN addresses a special case of VFL where the host has access only to labels and the (single) guest has access only to features. The original paper Gupta & Raskar (2018) also presents a protocol for training when the features are horizontally distributed across $N$ guests. Ceballos et al. (2020); Romanini et al. (2021) extend SplitNN for vertically partitioned data by vertically splitting the shallower layers (see Figure 1).
Figure 1: Diagram illustrating the distributed training of a neural network under SplitNN, VFL with SplitNN, and DVFL. The two input data for VFL and DVFL are partial features of the same entity.
Effectively, the guest models $m_{i}$ are ANNs and host model $m_{h}$ consists of a concatenation layer $m_{h}^{0}$ followed by an ANN $m_{H}^{1}$.
2.4 Impracticalities in VFL Training
Vertical Federated Learning is contingent on implicit assumptions regarding the reliability, trustworthiness and scale of participation.
Subversion of these assumptions can be catastrophic. A lost guest output or gradient is a single point of failure. The intersection $\mathcal{S}_{\mathrm{labels}}\cap\mathcal{S}_{\mathrm{guests}}$ may not be meaningfully large, especially when there are many guests. And, it is possible for a malicious guest to infer private data from feedback.
Thus far, the application of VFL has been limited to a small number of cross-silo settings where the silo owners are powerful enough to assume reliability and to afford the overhead of privacy guarantees. Further, cross-silo participation is small in number so scalability is not a concern. VFL is altogether infeasible for a large class of systems where devices are many or have low compute and communication bandwidths.
2.4.1 Single Points of Failure
In order to make a prediction on a batch of data, the host model $m_{h}:\mathbb{R}^{k_{1}}\times\dots\times\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ requires inputs from all guest models. If an input is unavailable to the host model, no prediction can be made– a catastrophic fault. Similarly, local models’ updates require gradients from the host, lacking which the model convergence slows down and may fail altogether. We consider the following fault model for a round of training:
  • Guest faults: A guest device may be unable to compute its output, update its model, or communicate.
  • Host faults: The host itself may fail, and be unable to receive inputs or calculate gradients.
  • Connection faults: The communication link between a guest and the host may be dropped during the forward or backward pass.
From a VFL host’s perspective, a guest fault and communication fault manifest as a missing input to its model $m_{h}(\text{ }\cdot\text{ };\theta_{h}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{\text{out}}$ during training. The host may employ one of two strategies:
  1. 1.
    Wait: The host may wait until all input arrives.
    • Under a guest fault or communication fault, the host will remain waiting indefinitely, and training will not continue.
    • If the host polls an unresponsive guest, it may receive an input intended for the next iteration, which will lead to incorrect model evaluation.
  2. 2.
    Timeout: The host may set a deadline for each round.
    • Under a guest fault or communication fault, training will fail since the host model is missing input.
    • If a message is successful but late, the model is still missing input.
    • If the model skips that training round and moves onto the next, it may receive the late message from the previous round, which will lead to incorrect model evaluation.
Therefore, a communication or guest fault leads to a catastrophic failure for training. When the host fails, the guest models are unable to update their parameters. Since the cardinality of the joint training dataset is that of the intersection of participants’ sample spaces $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$ (which may be small), losing updates from a single batch of data may be detrimental to model performance.
2.4.2 Data Intersection
Since VFL training is contingent on entity alignment, it is not usable in scenarios where the intersection of entities known to each guest is small. This is often the case when the number of guests is large. It could also be the case that, although $\mathcal{S}_{\mathrm{guests}}$ is reasonably large, labels are not available for all its members, i.e., $\mathcal{S}_{\mathrm{labels}}\cap\mathcal{S}_{\mathrm{guests}}\subset\mathcal{S}_{\mathrm{guests}}$.
2.4.3 Inference Attacks
Fu et al. (2022) demonstrate that it is possible for a malicious agent to directly infer the true class from gradients (e.g. by analyzing their sign) for many common loss functions during training. The authors also show that after VFL training is complete, it is possible for a curious guest to train its own prediction model if it can gain access to a very small number of labels. If a malicious guest inflates its learning rate during training, the aggregator gives more importance to results from its model. This way, a prediction model trained only on the features of the malicious guest is almost identical to the aggregator model itself. A malicious guest equipped with such a prediction head can make accurate guesses for labels without accessing other guests’ features.
Luo et al. (2021); Erdoğan et al. (2022) show that a generative network can be trained by a set of guests to model the distribution of other guests’ features. The authors also show that a malicious guest only needs to make a good guess for labels to directly infer features of other guests in vertically federated linear classifiers and decision trees.
Typical defenses include differential privacy, gradient compression, homomorphic encryption and secret shares for gradients and model features. However, these methods usually have a negative effect on model convergence rate and tightness Khan et al. (2023). Moreover, implementing such cryptographic algorithms is computationally expensive, and the cost increases with the number of participants.
3 Related work
3.1 Fault-Tolerant Distributed Optimization
The impact of redundancy for fault tolerance in distributed systems is well studied. Liu et al. (2021) provide a rigorous analysis of redundancy in the case of distributed optimization, showing that systems which are $2f$-redundant are approximately resilient to $f$ Byzantine agents. The authors also present two fault-tolerant algorithms for horizontally distributed convex optimization. Liu (2021); Bouhata et al. (2022) extensively review approaches to fault tolerance in distributed learning. Most methods require synchronization by an aggregator and incur computational overhead. Further, most methods involve manipulating gradients, which forces a choice between convergence rate and tightness.
For VFL: Few methods exist for to address fault tolerance in VFL. Li et al. (2023) introduce FedVS, an algorithm for systems where guest models are polynomial networks Livni et al. (2014a; b). FedVS uses a form of secret sharing to introduce redundancy in the input data, and can tolerate a fixed number of late or failed guests. However, FedVS and many of its predecessors Shi et al. (2022); Chen et al. (2020a); Hu et al. (2019) only address faults in the forward pass.
3.2 Federated Transfer Learning
Federated Transfer Learning (FTL) is a strategy for systems where there is some overlap in both sample space and feature space. The host owns labels and some features (i.e. it is an active participant) for entities in its sample space. Guests have limited labels for their task, but have some small overlap in feature space and sample space with the host (“aligned samples”).
Liu et al. (2020) present Secure Transfer Federated Learning (STFL), a framework for a single host and a single guest. Both parties learn a common latent space using local feature extractors. Encodings of the passive participant’s unaligned samples in this feature space can be used by the passive participant to learn its task model. Feng & Yu (2020); Feng et al. (2022) extend this idea to multiple parties.
For the related case of VFL with small $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$, Sun et al. (2023) introduce few-shot Vertical Federated Learning, where a few epochs of training on the aligned dataset is followed by semi-supervised learning on the non-intersecting entities of each client’s dataset. To generate labels, guests cluster the gradients they receive from the aligned samples. While this enables remarkable model performance, there is unwanted computational overhead required to generate the labels. Moreover, this is a form of label inference attack! Not only is this a breach of privacy to some degree, this clustering could be used downstream by a malicious agent to perform feature inference attacks (see Section 2.4.3).
3.3 Greedy and Localized Learning
Algorithms which update parameters in hidden layers of ANNs based on local or auxiliary objectives (as opposed to the global label loss) have recently seen renewed interest. In the scope of VFL training, they enable some useful properties for ANN training.
The primary drawbacks of BP in the scope of VFL are the inherent locking synchronization constraints, described in Jaderberg et al. (2017). In order to update its weights, a layer must wait for the backward signal to reach from the output layer (backward locking). In order to begin updating any layer, a full forward pass must first propagate to the end of the network (update locking). In order for a layer to evaluate an input to begin with, a forward pass must propagate to it (forward locking). These strict locks at the split layer are exactly the failure points for VFL training– forward locking causes training to fail if a guest does not communicate activations, backward locking causes training to fail if the host does not communicate gradients.
By eschewing the wait to evaluate the global objective, most greedy algorithms are backward unlocked if not update unlocked. In order to partially address forward locking for model parallel training Belilovsky et al. (2020) propose a replay buffer mechanism. Each agent responsible for training a module uses a replay buffer to store activations from the previous module’s agent. During a forward pass, the agent reads an activation from the buffer. If the forward signal has propagated through the previous module, it would have written the latest activation to the buffer. If not, the previous activation is reused.
4 Proposed method
DVFL splits the training into two phases coordinated by the owner. The first is an unsupervised and decoupled greedy guest-host representation learning phase. In this phase, the guest and host models learn feature extractors. The second is a supervised owner transfer-learning phase where the label owner learns to aggregate encodings from the host space and produce label predictions.
Within the guest-host phase, participants train their models asynchronously on their own obejctives. This allows some flexibility regarding when participants communicate and when they train. Since guests do not rely on feedback from hosts to update their models, host or communication faults are irrelevant to the guest models. Hosts store activations from guests as they arrive and use the latest message for their forward pass, so a missing input on the forward pass is no longer critical.
After hosts and guests complete training, the label owner trains its model. For each entity in $\mathcal{S}_{\mathrm{labels}}$, the owner requests guests and hosts to send its encoded and aggregated features. The owner uses this as an input for its task model, e.g. a classifier.
We present a DVFL approach to training split neural networks in the follwing subsections. The resultant algorithm is formalized in Algorithm 1. We also show how DVFL guests might exploit samples outside the joint sample space $\mathcal{S}_{\mathrm{guests}}$ to train better guest and host models.
4.1 The DVFL Hierarchy
Guests: The role of a guest in DVFL is identical to a VFL guest’s (see Section 2.1), i.e. guests $g_{i}\in\mathcal{G}$ are the agents $g_{i}$ who own datasets $\mathcal{D}_{i}$ with records of features $\mathcal{F}_{i}$ for the set of entities whose indices are present in $\mathcal{S}_{i}$. Each guest has a guest model and an unuspervised objective.
Hosts: Hosts are members of the set $\mathcal{H}$ of agents $h_{i}$ which own aggregating host models, which accept the concatenation of all guest model outputs as input. Each $h_{i}\in\mathcal{H}$ has:
  1. 1.
    Host model $m_{h,i}(\text{ }\cdot\text{ };\theta_{h,i}):\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}\times\dots\times\mathbb{R}^{k_{|\mathcal{G}|}}\to\mathbb{R}^{k_{h,i}}$ whose task is to aggregate and compress the data it receives from guests. It has its own unsupervised objective.
  2. 2.
    Single-reader single-writer “communication” registers $\mathbf{B}_{j,i}\forall j\in\{1,2,\dots,|\mathcal{G}|\}$. Each guest has a register on $h_{i}$, and writes its output to it. The host reads it to form its model input.
  3. 3.
    Input replay buffer $\mathbf{A}_{i}$. The host $h_{i}$ uses this to store a history of model inputs, if it wishes to reuse them.
Each host model has the same architecture but its own independent parameters $\theta_{h,i}$. It is possible that $\mathcal{H}\cap\mathcal{G}\neq\phi$.
Owner: The owner agent $o$ executes each training and communication round. It has access to ground truth labels, and is responsible for training a “final” or “owner” model on concatenated representations from members of $\mathcal{H}$ (i.e. it is a transfer learning model). The owner has the following:
  1. 1.
    Labels $L=\left\{\mathbf{y}_{j}\in\mathbb{R}^{\text{out}}:\text{label }\mathbf{y}_{j}\text{ of }\mathbf{x}_{j},j\in\mathcal{S}_{\mathrm{labels}}\right\}$:
    • $\mathcal{S}_{\mathrm{labels}}$ is the set of sample indices for which the host has the corresponding labels
    • We assume that the intersection of all the participants’ indices is not null to allow for training, i.e., $\left|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}\right|>0$.
  2. 2.
    Final model $m_{o}(\text{ }\cdot\text{ };\theta_{o}):\mathbb{R}^{k_{H}\cdot{|\mathcal{H}|}}\to\mathbb{R}^{\text{out}}$ with parameters $\theta_{o}$.
The label owner may be a member of $H\cup G$. In any case, we refer to it shorthand as “the owner”. Figure 2 illiustrates a DVFL system with three guests and two hosts.
Figure 2: A DVFL system with three guests and two hosts. Guests train their local models on unsupervised objectives and hosts also train their aggregating models on unsupervised objectives. After that, the label owner trains a transfer learning model (such as a linear classifier head) on the encodings from the hosts.
4.2 Asynchronicity in DVFL
To avoid a single point of failure, host models should continue training in the absence of inputs from some guests. We adapt the use of activation replay buffers of Belilovsky et al. (2020) to facilitate training under guest and communication faults. We use single-reader single-writer “communication” registers so that hosts can reuse old activations from failed guests. A host reads the latest messages in its registers and uses them to fulfill its model inputs.
If hosts wish to train for more iterations than they receive activations for, they may store their model inputs in replay buffers. They may then reuse inputs as needed. We discuss these mechanisms in further detail in Appendix D.
The task of a host’s model is only to learn an encoding of features in the joint guest domain, which is not an instance-specific task. Therefore, a host may still update its model meaningfully with some out-of-date activations.
4.3 Data beyond the intersection
Until now, we have assumed the input datasets to each model are “aligned”, viz. they only contain records for members of $\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$. The learning target of the guest models are to compress any data they receive as input, i.e. to encode features of a realization of guest domain’s data distribution. Similarly, each host learns to aggregate and compress encodings from all guests. These tasks depend on the input data’s domain (i.e. not on its instance). Therefore, guests may train their local models on features of $\mathbf{x}_{j}$ even when $j\in\mathcal{S}_{\mathrm{guests}}\setminus\mathcal{S}_{\mathrm{labels}}$.
4.4 DVFL for Split Training of Neural Networks
Given the DVFL hierarchy, ANNs are natural choices for participants’ models. Unsupervised neural networks are feature extractors Intrator (1992); Becker & Plumbley (1996), which is precisely the role of guest and host models. Similarly, the strong task performance of ANNs have been the primary driver of deep learning over the past few years, so they are suitable for owners, too. Indeed, just as SplitNN is a popular implementation of VFL, we expect ANNs to be common models for DVFL participants.
Figure 1 illustrates the relationship between DVFL and VFL in the context of a global ANN model. We may say $m_{i}$s are all ANNs with parameters $\theta_{i}$, $m_{h,j}$ are also ANNs whose input is the concatenation of the outputs of all $m_{i}$s, and $m_{o}$ is an ANN whose input is the concatenation of the outputs of all $m_{h,j}$s.
Algorithm 1 DVFL training for split neural networks
Input: Guest datasets $\mathcal{D}_{i}$, optimizers $\text{optim}_{g_{i}}$ $\forall i\in$$\{o\}\cup$$\{g_{1},\dots,g_{|\mathcal{G}|}\}\cup\{h_{1},\dots,h_{|\mathcal{H}|}\}$, communication schedule
Guest Training Round:
1:  for each guest $g_{i}$ in parallel do
2:     Take the next batch $\mathbf{x}_{t,\mathcal{F}_{i}}$ from dataset $\mathcal{D}_{i}$
3:     $\hat{\mathbf{x}}_{t,i}\leftarrow m_{i}(\mathbf{x}_{t,\mathcal{F}_{i}};\theta_{i})${Guest forward pass}
4:     if communication round then
5:        for each host $h_{j}\in\mathcal{H}$ do
6:           $\mathbf{B}_{i,j}\leftarrow\hat{\mathbf{x}}_{t,i}$
7:        end for
8:     end if
9:     Compute an unsupervised loss $\mathcal{L}_{i}(\hat{\mathbf{x}}_{t,i})$
10:     $\theta_{i}\leftarrow\text{optim}_{i}(\theta_{i},\nabla_{\theta_{i}}\mathcal{L}_{i})${Update guest model}
11:  end for
Host Training Round:
1:  for each host $h_{i}$ in parallel do
2:     if communication round then
3:        Read $\mathbf{B}_{j,i}\forall j\in\{1,2,\dots,|\mathcal{G}|\}$
4:        Concatenate all $\mathbf{B}_{j,i}$ and append to $\mathbf{A}_{i}$
5:     end if
6:     Read vector $\mathbf{b}_{t,i}$ from $\mathbf{A}_{i}$
7:     $\hat{\mathbf{b}}_{t,i}\leftarrow m_{h,i}(\mathbf{b}_{t,i};\theta_{h,i})${Host forward pass}
8:     Compute an unsupervised loss $\mathcal{L}_{h,i}\left(\hat{\mathbf{b}}_{t,i}\right)$
9:     $\theta_{h,i}\leftarrow\text{optim}_{h_{i}}(\theta_{h,i},\nabla_{\theta_{h,i}}\mathcal{L}_{h,i})${Update host model}
10:  end for
Owner executes:
1:  repeat
2:     Guest Training Round
3:     Host Training Round
4:  until convergence or a fixed number of iterations
5:  repeat
6:     for each entity $\mathbf{x_{j}}$, $j\in\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}$ do
7:        for all guests in parallel do
8:           Get features $\mathbf{x}_{j,\mathcal{F}_{i}}$ pertaining to entity $\mathbf{x_{j}}$
9:           Pass features through guest models and send activations to hosts
10:        end for
11:        for all hosts in parallel do
12:           Concatenate all received activations
13:           Pass activations through host models $m_{h,i}$ with parameters $\mathbf{\theta}_{h,i}$
14:           Send all activations to the owner
15:        end for
16:        Owner concatenates all inputs to form $\mathbf{x}_{j,\mathrm{enc}}$
17:        $\hat{\mathbf{y}}_{j}\leftarrow m_{o}({\mathbf{x}}_{j,\mathrm{enc}};\theta_{o})${Owner forward pass}
18:        Compute loss against labels $\mathcal{L}_{o}(\hat{\mathbf{y}}_{j},\mathbf{y}_{j})$
19:        $\theta_{i}\leftarrow\text{optim}_{o}(\theta_{o},\nabla_{\theta_{o}}\mathcal{L}_{o})${Update owner model}
20:     end for
21:  until convergence or a fixed number of iterations
5 Experiments
As a toy example, we consider a system of $|\mathcal{G}|$ guests which can each see $\frac{1}{|\mathcal{G}|}^{th}$ of an handwritten digit from the MNIST dataset. For our experiments, we set $|\mathcal{G}|=4$ i.e. $g_{1}$’s dataset contains the top $28\times 7$ pixel patch of an MNIST image, $g_{2}$’s dataset contains the second , $g_{3}$’s contains the third, and $g_{4}$’s contains the bottom $28\times 7$ patch (see Fig. 3).
Figure 3: For our experiments, the input received by each guest for a given sample, as well as the reconstruction its autoencoder model produces.
We use a learnable decoder head and MSE reconstruction as an unsupervised objective.
Baseline
DVFL is an alternate strategy to VFL. Naturally we compare DVFL for ANNs to SplitNN, a direct application of VFL to ANNs. Naive VFL implicitly employs the “wait” protocol discussed in Sec 2.4.1. There are some simple additions to SplitNN that would allow a more meaningful comparison. Firstly, if the host sets a deadline, we may continue training despite missing inputs by skipping (we call this SplitNN-skip) or by supplying replacement values. For replacement values, we try a zero tensor (SplitNN-zeros). Hosts may ask guests to send the identities of the entity to which each of their messages belong. This way, a late message cannot be believed to be intended for the next iteration. Results on more baselines and datasets are presented in Appendix A. The choice of baseline is discussed in Appendix C.
5.1 Fault Simulation
Six hyperparameters control the fault model described in Section 2.4.1. $R^{(d)}_{i}$ is the fault rate for a live $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host}\}$. $R^{(u)}_{i}$ is the chance a dead $i$ rejoins the network before the next request. We discuss how exactly the faults are simulated in Appendix B.2. We assume no faults during label supervision or inference.
5.2 Experimental Setup
Table 1: We compare DVFL to standard VFL. Here, we report the median and twice the standard deviation of MNIST test performance when trained with various entities $i\in\{\mathrm{connection,guest,host}\}$ susceptible to fail at a rate of $R^{(d)}_{i}=0.3$. A failed resource has $R^{(u)}_{i}=\mathbb{P}_{\mathrm{rejoin}}\in\{1,0.5,0.1\}$ to be available next time it is requested. SplitNN, the baseline VFL algorithm, cannot train at all when there are any faults. On the other hand, our DVFL implementation degrades gracefully with faults. Using an explicit fault handling strategy within VFL does allow it to train. But the degradation with faults is sharper than DVFL and the variance in model performance is higher. Under fault-free circumstances, SplitNN slightly outperforms our DVFL implementation. With minor adjustments, SplitNN can tolerate some faults, but our approach is more resilient and stable.
Strategy Test Accuracy (%) with 0.3 Loss Rate
No Faults Communication Loss Guest Loss Host Loss
$\mathbb{P}_{\mathrm{rejoin}}=$ N/A $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1 $\mathbb{P}_{\mathrm{rejoin}}=$ 1 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.5 $\mathbb{P}_{\mathrm{rejoin}}=$ 0.1
DVFL-NN 97.30$\pm$0.08 97.24$\pm$0.10 97.19$\pm$0.14 97.03$\pm$0.17 97.11$\pm$0.11 97.08$\pm$0.16 96.87$\pm$0.20 97.27$\pm$0.12 97.24$\pm$0.14 97.02$\pm$0.33
VFL (SplitNN) 97.80$\pm$0.17 N/A N/A N/A N/A N/A N/A N/A N/A N/A
VFL (SplitNN-skip) 97.84$\pm$0.16 97.50$\pm$0.54 96.60$\pm$0.51 77.21$\pm$5.34 97.44$\pm$0.57 96.65$\pm$0.45 80.22$\pm$4.31 97.65$\pm$0.41 97.59$\pm$0.50 97.00$\pm$0.89
VFL (SplitNN-zeros) 97.79$\pm$0.15 97.64$\pm$0.38 97.53$\pm$0.40 96.89$\pm$0.36 97.68$\pm$0.40 97.59$\pm$0.40 96.45$\pm$0.50 97.68$\pm$0.30 97.65$\pm$0.37 97.01$\pm$0.93
Performance under Faults
In order to show that it is possible to use DVFL even with crash faults, we train DVFL with $|\mathcal{H}|=4$ hosts under $R^{(d)}_{i}=0.3,R^{(d)}_{j}=0$ where $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host},\phi\},j\neq i$. Given a $R^{(d)}_{i}\neq 0$, we try $R^{(d)}_{i}\in\{0.1,0.5,1.0\}$.
Effect of Redundancy
In order to show that redundancy in hosts improves model performance and fault tolerance, we simulate DVFL training with $|\mathcal{H}|$ = 1, 2, 3, and 4 hosts. We measure test accuracy given $R_{\mathrm{connection}}\in\{0,0.3,0.6\}$ during training.
Learning from a Limited Intersection
We first fix a small intersection size, where all members of the intersection are labeled. We generate the remaining dataset for each device by shuffling the out-of-intersection samples and taking windows such that the remaining sample space is distributed evenly across all four guests. For example, if we wish to use $1024$ labeled samples, we shuffle samples indexed $1024-59999$ and split the $58976$ samples evenly, such that each guest gets $14744$ unique MNIST patches + $1024$ intersecting MNIST patches = $15768$ total samples. We try this experiment with $128$, $256$, $512$ and $1024$ labeled samples. We compare with SplitNN trained on only the labeled samples.
6 Results and Discussion
[…]
Figure 4: Model performance improves with redundancy, especially when there is a nonzero crash rate.
6.1 Graceful Degradation with Faults
We observe (Table 1) that model performance degrades with faults, but not catastrophically (unlike VFL with SplitNN). In ideal conditions, DVFL performance is slightly poorer than SplitNN. This is not unexpected, since our models are shallow. Wang et al. (2021) show performance improves with the length of backpropagation in blockwise-trained networks, and suggest it is due to increasing mutual information between shallower and deeper layers. Further, given a model, the reconstruction loss guides the feature extractor to features that will be useful for reconstruction– not necessarily good for classification. It is expected that the gap between DVFL and SplitNN will reduce with more general loss functions van den Oord et al. (2019); Chen et al. (2020b); Caron et al. (2020); Zbontar et al. (2021) and with self-supervised classification objectives. Further, observe that the degradation with faults in DVFL is much more graceful than our SplitNN variants. Not only does the SplitNN-wait/skip/stale performance drop more sharply, the variance increases more than it does with DVFL.
Guest faults are the worst
In the case of DVFL, the most severe type of fault is the guest loss. For example, when $R^{(d)}_{\mathrm{guest}}=0.3,R^{(u)}_{\mathrm{guest}}=0.1$ the accuracy drops from 97.30% to 96.87% while with communication and host loss it is 97.03 and 97.02% respectively. The performance of the combined model depends on unique outputs from guests– there is no redundancy. With a host or communication fault, guests are still able to update their own models (and there might be redundant hosts, too). In the case of SplitNN-skip and SplitNN-host, a communication fault is just as bad– a guest cannot send forward passes to (or receive backward passes from) the single host whether it or the connection is down. A host fault is equivalent to skipping that round for training.
$R^{(u)}=1$ simulates single-message loss, since the next time a message is sent or received, the participant is sure to have rejoined. $R^{(u)}=0.5$ simulates short bursts of loss. $R^{(u)}=0.1$ simulates appreciable downtime.
Tolerating some faults pays off
Induced by the artificial input, the Dropout effect Srivastava et al. (2014) leads to a reduction in bias. This competes against the loss of information in the model. In the case of DVFL, models do not accept feedback, so the scope for information loss is smaller. In the case of SplitNN with missing input imputation, models do not respond to faults as well as DVFL. Still, the test accuracy degrades somwewhat gracefully for this reason.
Other failure modes
In this study, we only examine loss faults. Other typical failure modes for distributed systems include arbitrary delays, incorrect order, and Byzantine faults. For the forward pass, the first two might be mitigated within VFL itself. Since hosts request encoded features from a particular sample or batch of samples, they may ask guests to send back the identity of those samples along with the features (as we assumed in the variants of SplitNN. While the privacy is dubious, there is precedent in VFL: see Sun et al. (2021); Yang et al. (2023); Jiang et al. (2022). But, in order to correctly serialize, the host must maintain a cache of unused messages. For the backwad pass, if a guest has sent two outputs before receiving a single gradient, it has no way of knowing which output it corresponds to. DVFL is immune to these issues, since guests and hosts train independently to each other.
6.2 Redundancy Improves Performance
We observe (Figure 4) that test accuracy improves with the number of hosts with all other system parameters fixed. This is not surprising, as the owner’s aggregation of host results can be seen as a form of bagging Breiman (1996). As a corollary, the degradation of performance with communication faults is more graceful with increasing redundancy.
The role of the DVFL host
Observe that host faults do not hurt DVFL performance very much. DVFL hosts’ primary function is aggregation. A DVFL host can achieve this without any learnable parameters, by simply concatenating inputs and passing them onto the owner when requested. The other role a DVFL host might play is to project the concatenated host output to a smaller dimension. What the host model must learn is to preserve useful features in the new dimension.
Table 2: We exploit data outside the sample intersection to train DVFL’s guests and hosts. For the owner model, we use labeled entities common to all datasets. This performs better than VFL, which can only use the common samples.
Strategy Test Accuracy (%)
# of Labeled Samples
128 256 512 1024
DVFL 76.45 82.51 85.18 88.90
VFL (SplitNN) 65.07 77.51 83.09 87.26
6.3 Out-of-Intersection Training
Although VFL with SplitNN outperforms our DVFL ANN algorithm in our experiments with perfect conditions, DVFL significantly outperforms VFL on a limited labeled intersection. DVFL’s advantage is greater with fewer labels (Table 2).
6.4 Security & Privacy
DVFL achieves these results while also completely eliminating the possibility of gradient-based inference attacks. As reviewed in 3, complex mechanisms can be developed on top of SplitNN to provide privacy guarantees within some bounds. DVFL achieves complete gradient privacy without any overhead or model convergence issues.
7 Conclusion and Future Work
We present Decoupled Vertical Federated Learning, a strategy for joint learning on vertically partitioned data. Instead of adding complexity to cope with the data leakage and synchronization associated with BP, we eschew feedback altogether. Combined with the ability to reuse feedforward signals, this allows for fault tolerance and privacy in training. Training may continue even when there are faults, or there are a limited number of samples common to all guests.
DVFL also provides a high degree of privacy. Malicious or compromised guests can no longer use gradient feedback to figure out possibly sensitive information. Existing techniques to enforce privacy guarantees either hamper model convergence, are expensive to implement or both.
Our results suggest its performance to be comparable to SplitNN on various machine learning tasks. Unlike SplitNN, DVFL degrades gracefully with various kinds of loss faults. Even with explicit fault handling strategies, SplitNN’s performance drops off more sharply than DVFL’s. In general, DVFL’s asynchronicity provides a great deal of engineering flexibility. Some of it is discussed in the Appendix.
The next step for DVFL is to generalize it for more tasks and general architectures. VFL is readily compatible with a variety of machine learning models (including ensemble models). With a judicious choice of training objective for guest models, DVFL may also be extended past ANNs.
Another possible direction for DVFL is the introduction of more elaborate fault handling strategies. In this work, the aggregating “host” uses the the latest activation received from a “guest” client. This allows for a simple implementation using single-reader single-writer registers. We observe that this leads to meaningful representation learning– empirically, we only observe a 0.26% drop in accuracy on MNIST even when there is a 30% crash fault rate. In a bid to close this gap further, hosts might build a model of the distribution of its inputs. Hosts could then sample missing inputs from its posterior conditioned on successful messages. A natural first step would be to use a running mean as an estimator.
Appendix A Results on More Datasets, Comparison to FedVS
In order to bolster the comparison to SplitNN, we evaluate the performance of DVFL on some benchmark vertically partitioned datasets. They are Credit Card Yeh & hui Lien (2009), Parkinson’s Sakar et al. (2019), CalTech-7 and Handwritten Markelle Kelly . Note that CalTech-7 is a subset of 7 classes CalTech 101 Li et al. (2022): Face, Motorbikes, Dolla-Bill, Garfield, Snoopy, Stop-Sign and Windsor-Chair. Both Handwritten and CalTech have 6 distinct “views”– features of distinct domains, e.g. spatial, frequency (Fourier), Gabor filter coefficients, etc.
FedVS Li et al. (2023) is a state-of-the-art VFL algorithm which aims to tackle both privacy and fault tolerance. There is a round of secret shares so that each guest holds an encrypted version of a horizontal section of other guests’ data and model parameters. The shares are arranged in such a manner that if $\leq L$ (a fixed system parameter) guests don’t send their outputs at the forward pass, there is enough redundant information in the other guests’ forward passes to account for it. In order to maintain privacy despite this sharing of data, guests quantize their models and data and apply randomly mask. This means there is no longer enough information for a malicious guest to exactly reconstruct the data it receives from other guests. But, this obscurity is at the cost of model performance. Further, this algorithm is contingent on the guests using polynomial networks Livni et al. (2014a; b).
In the following experiment, we compare our DVFL algorithm, SplitNN and FedVS. Our VFL and DVFL models are of the same depth as FedVS, but we use a smaller width. We also only use population normalization (as opposed to online batch and layer normalization). Therefore, the non-FedVS models have much fewer parameters and are simpler to train. We also take the liberty of using non-polynomial activations (Leaky ReLU). Model details and hyperparameters are in Appendix B. $|\mathcal{H}|=1$ for these experiments.
Table 3: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions. Even on tabular datasets harder (especially for ANNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN. Despite using a smaller model, both outperform the results cited in the FedVS paper Li et al. (2023).
Strategy Test Accuracy (%)
Dataset
Credit Card Parkinson’s CalTech-7 Handwritten
DVFL 82.66 85.76 99.00 98.25
VFL (SplitNN) 84.11 86.64 99.66 98.25
FedVS Li et al. (2023) [^2] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25
The feature splits and train-test splits for each dataset are identical to the FedVS paper Li et al. (2023). Further discussion about these results is available in Appendix C.
Appendix B Models, Hyperparameters and Implementation Details
B.1 Models
The “global”, single-participant neural network (see Figure 2) being split is a shallow multi-layer perceptron. It is parameterized by $W_{g}$ and $W_{h}$. $W_{g}$ is the dimension of the concatenated output of all guests viz. the input to a host. $W_{h}$ is the output dimension of a single host.
B.1.1 MNIST
The “guest” models have 2 layers, which squeeze the flattened $\frac{784}{|\mathcal{G}|}$-long input vector dimension to $\frac{400}{|\mathcal{G}|}$, and then to $\frac{W_{g}}{|\mathcal{G}|}$ with ReLU activation.
The “host” models squeeze the $W_{g}$-dimensional input to $\frac{W_{g}+3W_{h}}{4}$ and then to $W_{h}$, with Leaky ReLU activation (with negative half-plane slope = 0.01).
The “owner” model squeezes the $W_{h}\times|\mathcal{H}|$ dimensional input to $W_{h}$ and then to $40$ with a Leaky ReLU activation and finally to a $10$-class prediction with an implicit softmax via the cross entropy loss function.
For DVFL experiments, the guest, host, and owner modules are separate objects simulated to reside on separate devices. In order to produce a reconstruction, they also contain decoder MLPs which undo the squeezing operation. The guest decoder is activated by a sigmoid function (since the input image is normalized to the range $[0,1]$) and the host decoder is activated by a ReLU function (since the output of the guest model is also activated by a ReLU function and therefore cannot be negative).
For SplitNN experiments, the guest encoder architecture is identical to DVFL. The host model is a sequential module of the DVFL host encoder followed by the owner classifier.
We try DVFL for $W_{g}\in\{400,320,240,200\}$ and $W_{h}\in\{200,160,120\}$. The classification accuracies are presented in Table 4.
Table 4: Ablation study of model width (particularly, the dimensions of the outputs which are communicated) on model performance with DVFL. Our experiments in the main paper use $W_{g}=320$ and $W_{h}=160$.
$W_{h}$ Test Accuracy (%)
$W_{g}=200$ $W_{g}=240$ $W_{g}=320$ $W_{g}=400$
120 96.65 96.93 96.89 96.92
160 97.07 97.31 97.39 97.15
200 97.40 97.13 97.11 97.16
There two competing factors: the number of parameters (a monotonic function of $W_{g}\times W_{h}$) and the compression from host to guest. While more parameters implies a more generalizable (and hence hopefully accurate) model, an increase in $W_{g}$ for a fixed $W_{h}$ seems to suffer from some information loss, especially for larger $W_{h}$. It could be that the model eschews an encoding with more classification-relevant information when encodings with more reconstruction-relevant information are available thanks to the larger parameter space.
In the main paper, we use $W_{g}=320$ and $W_{h}=160$ for our experiments on MNIST. This choice enables good model performance while also demonstrating host models’ ability to compress the data.
B.1.2 VFL Datasets
Similar to MNIST, we use shallow multi-layer perceptrons for our results on vertically partitioned data in Appendix A. The model sizes for the guest $g_{i}$, host $h_{k}$ and owner $c$ are summarized in table 5, where $d_{i}$ is the input dimension of guest $g_{i}$.
Table 5: Model dimensions used in our experiments on tabular datasets. The dimension of the input to a guest $g_{i}$’s model is $d_{i}$. $W_{g}$ is the dimension of the concatenated output of the guests’ models. $W_{h}$ is the dimension of the output of one host’s model.
Dataset $W_{g}$ $W_{h}$ Activation
Guest Host
Credit Card $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ 10 Leaky ReLU Leaky ReLU
Parkinsons $\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor$ $\frac{\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor}{4}$ Leaky ReLU ReLU
CalTech-7 $256\times|\mathcal{G}|$ $256$ Leaky ReLU ReLU
Handwritten $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ $3\frac{\sum_{i}\lceil\frac{3d_{i}}{4}\rceil}{4}$ Leaky ReLU ReLU
B.2 Fault Simulation
We simulate faults based on six hyperparameters:
  1. 1.
    Communication Faults: During each communication round, for each pair $(g_{j},h_{i})$ we draw a sample from a uniform distribution with support $[0,1]$. We track the current status of each connection using flags.
    • $R^{(d)}_{\mathrm{connection}}$: If the connection between $(g_{j},h_{i})$ was alive on the last iteration, we kill it if the sampled value is greater than $R^{(u)}_{\mathrm{connection}}$.
    • $R^{(u)}_{\mathrm{connection}}$: If the connection between $(g_{j},h_{i})$ was dead on the last iteration, we revive it if the sampled value is lesser than $R^{(u)}_{\mathrm{connection}}$.
    In DVFL, we only write $g_{j}$’s activations to register $\mathbf{B}_{j,i}$ if. Similarly, for SplitNN, we check and update the flag before any forward pass from or backward pass to the split layer.
  2. 2.
    Guest Faults: Whenever a guest is called (i.e. there is a forward or backward pass), for each $g_{j}$ we draw a sample from a uniform distribution with support $[0,1]$.
    • $R^{(d)}_{\mathrm{guest}}$: If the guest $g_{j}$ was alive on the last call, we kill it if the sampled value is greater than $R^{(u)}_{\mathrm{guest}}$.
    • $R^{(u)}_{\mathrm{guest}}$: If the guest $g_{j}$ was dead on the last call, we revive it if the sampled value is lesser than $R^{(u)}_{\mathrm{guest}}$.
    In DVFL, we only update $\theta_{j}$ if the guest $g_{j}$ is alive. If it is a communication epoch, we only write $g_{j}$’s activations to $\mathbf{B}_{j,i}\forall i\in\{1,2,\dots,|\mathcal{H}|\}$ if the guest is alive. Similarly, in SplitNN, we only allow a guest to send a forward pass and receive gradients if it is alive.
  3. 3.
    Host Faults: During each training iteration, for each $h_{i}$ we draw a sample from a uniform distribution with support $[0,1]$.
    • $R^{(d)}_{\mathrm{host}}$: If the guest $h_{j}$ was alive on the last call, we kill it if the sampled value is greater than $R^{(u)}_{\mathrm{host}}$.
    • $R^{(u)}_{\mathrm{host}}$: If the guest $h_{j}$ was dead on the last call, we revive it if the sampled value is lesser than $R^{(u)}_{\mathrm{host}}$.
    In DVFL, we only update $\theta_{H,i}$ if the host is alive. If it is a communication epoch, we also drop the last item in input buffer $\mathbf{A}_{i}$. In the case of SplitNN, we assume the host is able to call on the guests before failing. When a SplitNN host fails, it does not compute any gradients.
B.3 Other Hyperparameters
For all experiments, we perform a hyperparameter search. The values we used for the results presented in this paper are below.
B.3.1 MNIST
For guests and DVFL hosts, we use an Adam optimizer with the default $\beta_{1}=0.9,\beta_{2}=0.999$. We use a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$. For DVFL owners and SplitNN guests we use stochastic gradient descent with a learning rate of $10^{-2}$ and momentum of $0.5$. In general, guests were trained for 20 epochs, hosts for 40 and owners for 60. For SplitNN, the whole model was trained end-to-end with 60 epochs. For the experiments on limited intersecting datasets, owners/SplitNN hosts are trained for 160 epochs. In all cases, DVFL owner models and SplitNN host models were trained with early stopping.
B.3.2 VFL Datasets
Hyperparameters are presented in Table 6.
Table 6: Hyperparameters for the results presented in Appendix A.
Dataset Learning Rate
DVFL SplitNN
Guest Host Owner Guest Host
(Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) (Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) (SGD, momentum = 0.5) (Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999) (Adam, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999)
Handwritten 2e-3 1e-3 3e-3 1e-3 1e-4
CalTech-7 5e-4 1e-2 3e-5 1e-4 1e-4
Parkinson’s 5e-3 4e-3 5e-4 1e-4 5e-5
Credit Card 1e-3 1e-3 1e-3 1e-3 1e-3
B.4 Implementation Details
We simulate parallel training by time multiplexing and using preallocated replay and communication tensors. The experiments were run on a variety of hardware, including mobile CPUs (Apple M2) and workstation GPUs (on a single NVIDIA RTX A5000 and single-GPU, and on a single NVIDIA Tesla K40).
Appendix C Choice of Baseline
C.1 SplitNN as a Baseline
In this work we present DVFL, an alternate strategy to VFL. We demonstrate its prowess through a DVFL implementation for ANNs. Since DVFL is altogether a new strategy to train on vertically partitioned data, it makes sense to benchmark it against the current status quo for VFL with ANNs, which is SplitNN Gupta & Raskar (2018); Ceballos et al. (2020).
SplitNN with a timeout
In its base form, the SplitNN algorithm does not have any explicit fault handling strategy. Implicitly (assuming that the coordination between hosts and guests is interrupt-driven), the hosts use the “wait” strategy. Training stops completely if inputs are missing. We also consider strategies using “timeout”. Timeout allows training to continue, although hampered. We handle timeouts by skipping or replacing the values. If a message appears after the deadline, it needs to be discarded lest it appear as if it were intended for the next iteration. This artificial baseline represents a best-case fault-tolerance for vanilla VFL. This way, it gets a fair shot against DVFL. Indeed, since DVFL suffers from some information loss (since feedback is very limited), VFL equipped with these strategies may still outperform DVFL at low fault rates.
Busy-waiting SplitNN
Under “wait”, training can be made to restart if the failed participant comes back online by busy-waiting viz. the repeatedly polling a failed guest. If the guests send sample IDs along with their features (and hosts cache messages until they are used), they may figure out which iteration a particular message is intended for. This way, no information is lost. However, until that failed participant is back online, training is held up. This is also a poor strategy from a systems perspective, since it wastes bandwidth and computational resources, especially when there is a message authentication in place. If a guest is laggy and the host polls it many times before it repsonds, the cache of messages the host must store will be very large. Neither in this case nor the timeout case is there a well-defined strategy to handle backward-pass faults.
C.2 Other Possible Baselines
DVFL addresses various VFL issues such as fault tolerance, privacy and sample size with one fell swoop. There are various mechanisms built upon standard VFL algorithms (including SplitNN) which aim to address some or the other of these issues. FedVS Li et al. (2023) is a work that addresses multiple, i.e. fault tolerance and privacy. FedVS might be a reasonable benchmark for DVFL, and we have presented some results against its paper in Appendix A. DVFL outperforms FedVS in perfect conditions. In the main paper, we showed that DVFL performance does not degrade much with faults with our generalized fault model. On the other hand, FedVS performance is completely fault tolerant, except under a much more stringent fault model. Beyond this, FedVS constrains guest models to use a specific type of model– polynomial networks Livni et al. (2014a; b). Moreover, FedVS only deals with faults during the forward pass. FedVS considers a restricted fault model: upto a fixed number of guests can be late to/fail to communicate their activations with the hosts. These are also true for most of its predecessors e.g. Chen et al. (2020a); Hu et al. (2019); Shi et al. (2022). DVFL does not require any such system parameter, and is tolerant to a broader class of faults where any participant can fail at any time, including during backpropagation.
Appendix D Communication and Space Complexity
D.1 Periodic Communication
The owner initiates every round of training as well as communication for all agents. Since the training of guests and hosts are decoupled, the owner may ask guests to communicate their latent activations (and therefore start host training rounds) at arbitrary intervals. In VFL, guest model updates are dependent on host feedback. So there is every-shot communication, i.e. guests send every single output to the host.
With one-shot communication such as FedOnce Wu et al. (2022), the volume of communication is greatly reduced, at the cost of model performance. Equivalently, Belilovsky et al. (2020); Siddiqui et al. (2023) demonstrate that updating all layers on local objectives on each forward pass performs better than fully training and freezing a layer before moving on to the next.
We study the tradeoff between communication cost and model performance by introducing a hyperparameter $K$, which is the period (in guest epochs) at which communication epochs occur. During a communication epoch, every iteration of the guest model iteration is followed by a communication round i.e. guests write their activations to the communication registers, and hosts read them, concatenate them and store the concatenations in their input buffers.
If there is a host training iteration at every guest training iteration communication round, after the guests complete $N$ epochs, the hosts will have
  1. 1.
    A model trained for $\left\lfloor\frac{N}{K}\right\rfloor$ epochs
  2. 2.
    The history of model inputs from each iteration in $\left\lfloor\frac{N}{K}\right\rfloor$ epochs of training
If the host must complete a certain number $M>\frac{|\mathcal{S}_{\mathrm{guests}}|}{\text{batch size}}\times\left\lfloor\frac{N}{K}\right\rfloor$ iterations of training, it reuses these activations until the required number of iterations is met.
We simulate VFL training with $4$ hosts and 00 fault rate. We measure test accuracy for the following communication periods:
  • Every guest epoch ($K=1$)
  • Every 5 guest epochs ($K=5$)
  • Every 10 guest epochs ($K=10$)
  • Only once, after 20 guest epochs ($K=20$).
Note that the dimension of the latent representations communicated by each guest to reach host is $80$, and each dimension corresponds to a 32-bit floating point number. The MNIST training dataset has $60000$ samples. If the total number of guest training epochs is $N$ and the period of communication epochs is $K$, the total communication cost per guest is ||×$|\mathcal{H}|\times\lfloor\frac{N}{K}\rfloor\times 60000\times 80\times 32$ bits. Each message received is saved by the host in the activation replay [^3] Although this is not necessary if the host will not reuse activations, viz. a host wants to train for $k$ iterations and it receives activations of $k$ iterations from the guests. This is discussed in more detail in Appendix D.2.
0055551010101015151515202020202525252596.996.996.996.99797979797.197.197.197.197.297.297.297.297.397.397.397.397.497.497.497.4Number of HostsMNIST Test Accuracy (%)
Figure 5: Model performance degrades with an increase in communication period, i.e. more communication is correlated with better performance.
Consistent with results from Belilovsky et al. (2020); Wu et al. (2022); Siddiqui et al. (2023), training the depth-wise split model parallely yields better model performance. We observe the tradeoff between communication cost and model performance. With 1.536GB, 0.307GB, 0.154GB and 0.077GB outgoing from each guest, the accuracies on the MNIST test set are $97.30\%$, $97.14\%$ $97.08\%$ and $96.84\%$ respectively.
D.2 Activation Replay Mechanism & its Computational Overhead
A host reads its model inputs from its activation replay buffer. Thanks to this, it can train asynchronously to the guests. The host can train independently to the guests as long as the guests have communicated at some prior point of time. In fact, that is the case in our experiments– we simulate host models’ training offline, after guest training. As discussed in Appendix D.1, this allows control over when the guests communicate.
Another way to look at this is: hosts do not need to train for the same number of iterations as guests communicate messages for. This includes the case where guests communicate every iteration, but hosts want to train more (or less) iterations. As we detailed in Appendix B.3, we actually implement this for our experiments. While we use 40 epochs to train the entire SplitNN model, we only use 20 epochs to train the DVFL guests. Guests need not remain online for any more time than required to train 20 epochs. So, not only does this allow for more flexibility, it also helps fault tolerance.
Computational Resources
In order to maintain this asynchronicity, hosts need to maintain the activation replay buffer. For each message it receives, the size of the buffer grows by batch $\text{batch size}\times W_{g}$. Further, these expensive memory read/write operations also add time complexity to DVFL hosts’ training. This overhead does not exist in standard VFL. This it is DVFL’s primary drawback [^4] The other source of overhead is the calculation of the self-supervised/unsupervised loss. For example, consider our experiments with MSE as the guest and host objectives. Along with our model, we need to train a decoder head that projects the guest output onto a space of the same dimension as input. Modern contrastive self-supervised losses e.g. Chen et al. (2020b); Zbontar et al. (2021) typically use a shallow projector network from the desired representation onto a latent space. The loss is calculated using the projected vectors. Intuitively, in the absence of a pre-computed ground truth label, the model must perform extra computation to generate the “self-”supervisory signals., although it can be mitigated to some extent since some models can be trained for fewer epochs as discussed above. However, it is important to note that it is not strictly necessary to save all activations. If a host trains for $\leq$ the number of epochs that guests train/communicate for, it does not need to reuse activations. In that case, a host model input can be deleted after it is used once.
[…]